# Higgs Phases and Boundary Criticality in memoriam Peter Higgs

### v1

@@ III.4 Boundary Symmetry Breaking in Non-Abelian Fundamental-Higgs Models
The boundary action is given by Eq. 42. For $N>2$ the Higgs rotors can no longer be identified as group elements. Because of this, the $\cramped{\kappa\to\infty}$ constraint,
@@
does not completely trivialize the bulk. To see this, note that the constraint enforces that nearest-neighbor Higgs rotors are parallel relative to the gauge field. However, since there is still freedom to perform rotations about the colinear axis of the remaining $N-1$ components of the Higgs field, the gauge field can continue to fluctuate so long as it does not rotate the Higgs field away from the parallel axis. This is explicitly seen by rotation to unitary gauge where the Higgs field is parallel in a global frame, which then fixes one diagonal element of the gauge field to unity, i.e.
[…]
@@
[…]
Thus in this limit the bulk theory becomes (gauge-equivalent to) an $\mathrm{SU}(N-1)$ gauge theory. Note that $\mathrm{SU}(1)$ is trivial, making the $\mathrm{SU}(2)$ case special, as discussed in Section III.2.
III.4.1 Gauge-Invariant Resolution of the Infinite Kappa Limit
To resolve the constraint in a fully gauge-invariant fashion analogous to Eqs. 22 and 60, we first note that the constraint Eq. 62 actually implies that
@@ III.4.1 Gauge-Invariant Resolution of the Infinite Kappa Limit
While Eq. 66 demonstrates the reduction of the gauge group, it is not that useful for formulating the boundary theory. A more useful way is to decompose each $U_{ij}$ by sandwiching it between two resolutions of the identity decomposed into the parallel and perpendicular subspace of $|\phi_{i}\rangle$ and $|\phi_{j}\rangle$. Namely, for site $i$
@@
In the limit $\kappa\to\infty$, this simplifies significantly. Firstly, the two cross terms (the last line) are exactly zero by Eq. 66, i.e. because $\cramped{P_{i}U_{ij}|\phi_{j}\rangle=P_{i}|\phi_{i}\rangle=0}$. Second, in the second term, Eq. 65 reduces it to $|\phi_{i}\rangle\langle\phi_{j}|$. In summary, in the limit $\kappa\to\infty$ every link variable in the bulk can be expressed as
@@
Now consider the three-legged plaquettes appearing in the boundary action, Eq. 42. Inserting the identities at the two bulk sites, we obtain
@@
where we have defined the gauge-invariant variables
$|\Phi_{i}\rangle\equiv U_{i}|\phi_{i}\rangle,$ (72)
which are $\mathbbm{C}^{N}$ vectors corresponding to the short half-open Wilson strings at the boundary. This highlights an important distinction for fundamental Higgs compared to group-valued Higgs models—here the half-open Wilson string is a vector degree of freedom, not group-valued. The $\kappa\to\infty$ theory then can be expressed in the gauge-invariant form
[…]
Thus the bulk Higgses down to an $\mathrm{SU}(N-1)$ gauge theory while the boundary decomposes into a pure $\mathrm{SU}(N-1)$ part plus a part which may be viewed as an $\mathrm{SU}(N)$ ferromagnet. The global $\mathrm{SU}(N)$ boundary symmetry acts as $U_{i}\to gU_{i}$ and simultaneously $|\Phi_{i}\rangle\to g|\Phi_{i}\rangle$. We generically expect this $\mathrm{SU}(N)$ symmetry to spontaneously break above a critical $\beta$ down to $\mathrm{SU}(N-1)$, and the short Wilson string rotors $|\Phi_{i}\rangle$ on the boundary to exhibit long range order and Goldstone modes. We discuss the nature of this phase transition further in Section III.5.
Figure 8: Location of the boundary phase transition for the SU(2) Higgs model for different values of $\alpha=\beta_{\text{bdry}}/\beta_{\text{bulk}}$, determined by the location of the peak in the boundary order parameter susceptibility (c.f. Fig. 7) . Large values of $\alpha$ push the transition line to smaller $\beta_{\text{bulk}}$, while small values of $\alpha$ push the phase boundary to larger $\beta_{\text{bulk}}$. The cyan square indicates the location of the bulk critical endpoint. For $\alpha=1$ (red), the boundary transition line appears to terminate at the bulk critical point, and closely follows the bulk rapid-crossover (super critical) region that extends beyond the critical endpoint. This remains true for $\alpha>1$, while for sufficiently small $\alpha<1$ the boundary transition line appears to terminate on the bulk first-order transition line.
III.4.2 Formulation for General Gauge Groups
@@
Assuming the bulk is gapped with a finite correlation length, as it must be if $\mathcal{H}$ is non-Abelian, the boundary is quasi-$(D-1)$-dimensional with some finite correlation length extending into the bulk. The system retains a $\mathcal{G}$ global symmetry rotating all of the boundary links, together with the bulk $\mathcal{H}$ gauge symmetry. The appropriate boundary theory is therefore expected to be a gauged nonlinear $\sigma$-model with target space $\mathcal{G}$ with subgroup $\mathcal{H}$ gauged, or equivalently, a nonlinear $\sigma$-model with target space the quotient space $\mathcal{G}/\mathcal{H}$. For example, $\mathrm{SU}(N)$ Higgses down to $\mathrm{SU}(N-1)$ and the quotient space is $\cramped{\mathrm{SU}(N-1)/\mathrm{SU}(N)\simeq S^{2N-1}}$, which agrees with the finding in Eq. 73 of a boundary theory of $\mathrm{SU}(N)$ rotors, whose configuration space is a sphere in $2N$ dimensions. Similarly, $\mathrm{SO}(N)$ Higgses to $\mathrm{SO}(N-1)$, with quotient space $S^{N-1}$.
III.4.3 Hamiltonian Perspective
Similar considerations apply to the Hamiltonian formulation of the non-Abelian Higgs theory described in Section III.1.2. The analog of the large $\kappa$ limit in the Hamiltonian formulation (Eq. 33) is first of all to drop the conjugate variables $\hat{Q}_{i}^{m}$ leaving only the $\hat{\Lambda}$ variables in the matter sector. This essentially renders the Higgs fields classical and they may be gauge fixed without loss of generality along some fixed direction $\phi^{\alpha}=\delta^{\alpha 1}$. There is now a Hamiltonian constraint
$-\kappa\sum_{\ell}\hat{U}_{\ell}^{11}$ (75)
that breaks gauge fluctuations from $SU(N)$ down to $SU(N-1)$.
Now if we consider only the boundary plaquettes, this constraint acts only on the bulk links parallel to the boundary while those extending out of the boundary have no such constraint. Therefore the boundary theory of the four dimensional bulk in the large $\kappa$ limit is a three dimensional $SU(N)$ chiral model that is partially gauged by an $SU(N-1)$ gauge group where the $SU(N-1)$ gauge theory permeates the bulk.
III.5 Boundary Phase Transition Order and Universality Class
So far we have demonstrated that Higgs models with group-valued or fundamental Higgs fields have large-$\kappa$ limits with well-defined boundary degrees of freedom that may exhibit symmetry breaking. Having discussed the $\kappa\rightarrow\infty$ boundary actions for the family of fundamental Higgs SU($N$) models we are in a position to say something about the phase transitions at the boundary. As the bulk is gapped for finite $\kappa$ we should expect that the boundary theory has a finite correlation length into the bulk making it a quasi-($\cramped{D-1}$)-dimensional boundary theory so that statements at $\kappa\rightarrow\infty$ hold also for finite $\kappa$.
@@ III.5 Boundary Phase Transition Order and Universality Class
III.5.1 Group-Valued Higgs
The boundary degrees of freedom, Eq. 59, are gauge-invariant composites transforming under the global $\mathcal{G}_{L}\times\mathcal{G}_{R}$ symmetry. Considering the case $\mathcal{G}=\mathrm{SU}(N)$, the coarse-grained (generally complex) matrix-valued fields are denoted $X^{ab}_{i}$ [40], where $a,b$ are color indices, which transform as $\mathbf{X}\rightarrow\mathbf{g}_{L}\mathbf{X}\mathbf{g}_{R}^{\dagger}$ under the symmetry. The Landau theory is
@@ III.5.1 Group-Valued Higgs
Chiral models on a lattice, such as the one in Eq. 60, have been studied for many years especially in two dimensions at large $N$, where they are integrable [41]. If one is interested in boundaries of four dimensional gauge theories, the three dimensional analogs of such models are of interest. One early work on $\mathcal{G}=SU(N)$ in three dimensions, Ref. [40], contains Monte Carlo results for $N=2,3,4$. For $N=2$ (Section III.2) the symmetry group is $\cramped{\mathrm{SU}(2)\times\mathrm{SU}(2)\simeq\mathrm{O}(4)}$ and the transition is continuous, consistent with O($4$) criticality. For $N=3,4$ the numerical results reveal the transition to be first order in agreement with the mean field theory predictions.
@@
III.5.2 Fundamental Higgs and SU(3) Numerical Results
The boundary model in the case of a fundamental Higgs has gauge-invariant degrees of freedom of the form $\Phi^{a}_{i}\equiv U_{i}^{ab}\phi_{i}^{b}$ where $a,b$ are the color indices. There is a single global $\mathcal{G}$ symmetry that acts from the left $\Phi\rightarrow g\Phi$. This will be broken spontaneously for sufficiently large $\beta$ (modulo Mermin-Wagner restrictions). The coarse-grained version of $\Phi^{a}$ is denoted $\Psi^{a}$ and the Landau theory for SU($N$) is
which is invariant under U($N$) transformations. This may instead be viewed as a theory of $2N$ real variables invariant under the enlarged O($2N$) symmetry group. This Landau theory gives the impression that, for the three dimensional boundary of a four dimensional SU($N$) gauge theory, there is a phase transition in the O($2N$) universality class. In the case of SU($2$) this is O($4$) criticality as shown above at the level of the microscopic model both analytically and numerically. Note that the coupling of the $\Phi_{i}$ rotors in Eq. 73 is invariant under O($2N$) rotations, even though the microscopic action manifestly only has a global $\mathrm{SU}(N)$ symmetry.
It may seem surprising that the SU($N$) invariant model exhibits O($2N$) criticality. One simple check is that SU($N$) has $N^{2}-1$ generators and that SU($N$$\rightarrow$ SU($N-1$) symmetry breaking therefore has $12N-$ 1 broken generators (corresponding to the number of Goldstone modes) which matches the count for O($2N$$\rightarrow$ O($2N-1$) symmetry breaking where O($N$) has $N(N-1)/2$ generators. But what about the terms that are SU($N$) invariant but not O($2N$) invariant? These terms are certainly present. An example is given by taking an operator $\Xi_{ab}\equiv\Psi_{a}\Psi^{\dagger}_{b}$ which transforms as $g\Xi g^{\dagger}$ and considering its determinant. This is only U($N$) invariant, but it is also an irrelevant operator for $N>2$ in $D=4$. More precisely, the couplings in the action originating from the determinant have mass dimension $[g_{\rm det}]=D+N(2-D)$ which, in dimensions higher than two, is negative for all but $N=2,3$ in three dimensions and $N=2$ in four dimensions. As we have seen, the case of SU($2$) is special because it is identical to a problem with a group valued Higgs. So we have treated it separately. Therefore the remaining puzzle relates to SU($3$) in three dimensions where the determinant coupling goes like $|\Psi^{\dagger}\Psi|^{3}$ and is marginal by power counting. As with the problem of scalar field theory in three dimensions [43], among other cases, we expect this sixth order term to be marginally irrelevant. Taking this together with our results for SU($2$) we surmise that the SU($N$) boundary theory phase transition is in the O($2N$) universality class since terms breaking O($2N$) down to SU($N$) are irrelevant or marginally irrelevant.
The discussion above shows that we should set aside the three dimensional boundary theory of the SU($3$) Higgs model in four dimensions for further examination. We anticipate a boundary phase transition in the O($6$) universality class for this case. This model is also interesting because, unlike the SU($2$) Higgs model, it has a dynamical bulk as the Higgs mechanism in this case leaves residual bulk degrees of freedom. We have therefore performed simulations of SU($3$) gauge theory with a fundamental Higgs field. The Higgs field $\phi_{i}$ is a $3$-component complex vector at each site, and we define the average of the Higgs field measured from the vacuum end of the boundary links,
$\mathcal{O}=\frac{1}{L^{3}}\sum_{i\in\partial X}(U_{i}\phi_{i})\,\in\,\mathbbm{C}^{3}$ (78)
The numerical results are presented in Fig. 9. Figure 9(a) shows the behavior of this order parameter for various system sizes, while Fig. 9(b) shows the behavior of its variance (susceptibility). The behavior is qualitatively consistent with an order parameter at a second order phase transition, approaching zero at small $\beta$ and continuously increasing from $\beta>\beta_{c}\approx 4.19$, with a diverging susceptibility. Figure 9 shows the Binder cumulant scaled by the 3D O($6$) critical exponent $\nu\approx 0.789$ [42], which shows a clear collapse to a universal scaling function. We therefore conclude that the SU($3$) Higgs model in $D=4$ indeed demonstrates a boundary phase transition in the $3$D O($6$) universality class, even though a single fundamental Higgs field does not remove all bulk degrees of freedom when $\kappa\to\infty$, in agreement to the general picture of boundary O($2N$) criticality on the boundary of $\mathrm{SU}(N)$ fundamental Higgs models.
III.6 Summary and Discussion: Non-Abelian Case

### v2

@@ 3.4 Boundary Symmetry Breaking in Non-Abelian Fundamental-Higgs Models
The boundary action is given by Eq.˜42. For $N>2$ the Higgs rotors can no longer be identified as group elements. Because of this, the $\cramped{\kappa\to\infty}$ constraint,
@@
does not completely trivialize the bulk. To see this, note that the constraint enforces that nearest-neighbor Higgs rotors are parallel relative to the gauge field. However, since there is still freedom to perform rotations about the colinear axis of the remaining N1N-1 components of the Higgs field, the gauge field can continue to fluctuate so long as it does not rotate the Higgs field away from the parallel axis. This is explicitly seen by rotation to unitary gauge where the Higgs field is parallel in a global frame, which then fixes one diagonal element of the gauge field to unity, i.e.
$\phi_{i}=\begin{pmatrix}1\\ 0\\ \vdots\end{pmatrix}\quad\xRightarrow{\kappa\to\infty}\quad(U_{ij})^{11}=1\qquad\text{(unitary gauge)}.$ (63)
@@
$U_{ij}^{\mathrm{SU}(N)}\xrightarrow{\kappa\to\infty}\begin{pmatrix}1&0\\ 0&U^{\text{SU}(N-1)}_{ij}\end{pmatrix}\qquad\text{(unitary gauge)}.$ (64)
Thus in this limit the bulk theory becomes (gauge-equivalent to) an $\mathrm{SU}(N-1)$ gauge theory. Note that $\mathrm{SU}(1)$ is trivial, making the $\mathrm{SU}(2)$ case special, as discussed in Section˜3.2.
3.4.1 Gauge-Invariant Resolution of the Infinite Kappa Limit
To resolve the constraint in a fully gauge-invariant fashion analogous to Eqs.˜22 and 60, we first note that the constraint Eq.˜62 actually implies that
@@ 3.4.1 Gauge-Invariant Resolution of the Infinite Kappa Limit
While Eq.˜66 demonstrates the reduction of the gauge group, it is not that useful for formulating the boundary theory. A more useful way is to decompose each $U_{ij}$ by sandwiching it between two resolutions of the identity decomposed into the parallel and perpendicular subspace of $|\phi_{i}\rangle$ and $|\phi_{j}\rangle$. Namely, for site $i$
@@
In the limit $\kappa\to\infty$, this simplifies significantly. Firstly, the two cross terms (the last line) are exactly zero by Eq.˜66, i.e. because $\cramped{P_{i}U_{ij}|\phi_{j}\rangle=P_{i}|\phi_{i}\rangle=0}$. Second, in the second term, Eq.˜65 reduces it to $|\phi_{i}\rangle\langle\phi_{j}|$. In summary, in the limit $\kappa\to\infty$ every link variable in the bulk can be expressed as
@@
Now consider the three-legged plaquettes appearing in the boundary action, Eq.˜42. Inserting the identities at the two bulk sites, we obtain
$\Tr[U_{i}U_{ij}U_{j}^{-1}]\xrightarrow{\kappa\to\infty} Tr$$\Tr[U_{i}\left(P_{i}U_{ij}P_{j}+|\phi_{i}\rangle\langle\phi_{j}|\right)U_{j}^{-1}]$
$=\,$ Tr[$\Tr[U_{i}P_{i}U_{ij}P_{j}U_{j}^{-1}]+\langle\Phi_{i}|\Phi_{j}\rangle$ (72)
where we have defined the gauge-invariant variables
$|\Phi_{i}\rangle\equiv U_{i}|\phi_{i}\rangle,$ (73)
which are $\mathbbm{C}^{N}$ unit vectors corresponding to the short half-open Wilson strings at the boundary. This highlights an important distinction for fundamental Higgs compared to group-valued Higgs models—here the half-open Wilson string is a vector degree of freedom, not group-valued. The $\kappa\to\infty$ theory then can be expressed in the gauge-invariant form
@@
Thus the bulk Higgses down to an $\mathrm{SU}(N-1)$ gauge theory while the boundary decomposes into a pure $\mathrm{SU}(N-1)$ part plus a part which may be viewed as an $\mathrm{SU}(N)$ ferromagnet. The global $\mathrm{SU}(N)$ boundary symmetry acts as $U_{i}\to gU_{i}$, under which $|\Phi_{i}\rangle\to g|\Phi_{i}\rangle$. The three-leg projected Wilson loop which appears in the boundary action in Eq.˜74 is not charged under this symmetry since the trace cancels the contribution from the two ends. We therefore generically expect the boundary $\mathrm{SU}(N)$ symmetry to spontaneously break above a critical $\beta$ down to $\mathrm{SU}(N-1)$, and the short Wilson string rotors $|\Phi_{i}\rangle$ on the boundary to exhibit long range order and Goldstone modes. We discuss the nature of this phase transition further and provide numerical support for this statement in Section˜3.5.
3.4.2 Formulation for General Gauge Groups
@@
Assuming the bulk is gapped with a finite correlation length, as it must be if $\mathcal{H}$ is non-Abelian, the boundary is quasi-(D1)(D-1)-dimensional with some finite correlation length extending into the bulk. The system retains a $\mathcal{G}$ global symmetry rotating all of the boundary links, together with the bulk $\mathcal{H}$ gauge symmetry. The appropriate boundary theory is therefore expected to be a gauged nonlinear $\sigma$-model with target space $\mathcal{G}$ with subgroup $\mathcal{H}$ gauged, or equivalently, a nonlinear $\sigma$-model with target space the quotient space $\mathcal{G}/\mathcal{H}$. For example, $\mathrm{SU}(N)$ Higgses down to $\mathrm{SU}(N-1)$ and the quotient space is $\cramped{\mathrm{SU}(N-1)/\mathrm{SU}(N)\simeq S^{2N-1}}$, which agrees with the finding in Eq.˜74 of a boundary theory of $\mathrm{SU}(N)$ rotors, whose configuration space is a sphere in $2N$ dimensions. Similarly, $\mathrm{SO}(N)$ Higgses to $\mathrm{SO}(N-1)$, with quotient space $S^{N-1}$.
3.4.3 Hamiltonian Perspective
Similar considerations apply to the Hamiltonian formulation of the non-Abelian Higgs theory described in Section˜3.1.2. The analog of the large $\kappa$ limit in the Hamiltonian formulation (Eq.˜33) is first of all to drop the conjugate variables $\hat{Q}_{i}^{m}$ leaving only the $\hat{\Lambda}$ variables in the matter sector. This essentially renders the Higgs fields classical and they may be gauge fixed without loss of generality along some fixed direction $\phi^{\alpha}=\delta^{\alpha 1}$. There is now a Hamiltonian constraint
$\kappa\sum_{\ell}\hat{U}_{\ell}^{11}$ (76)
that breaks gauge fluctuations from $SU(N)$ down to SU(N1)SU(N-1).
Now if we consider only the boundary plaquettes, this constraint acts only on the bulk links parallel to the boundary while those extending out of the boundary have no such constraint. Therefore the boundary theory of the four dimensional bulk in the large $\kappa$ limit is a three dimensional $SU(N)$ chiral model that is partially gauged by an SU(N1)SU(N-1) gauge group where the SU(N1)SU(N-1) gauge theory permeates the bulk.
3.5 Boundary Phase Transition Order and Universality Class
So far we have demonstrated that Higgs models with group-valued or fundamental Higgs fields have large-$\kappa$ limits with well-defined boundary degrees of freedom that may exhibit symmetry breaking. Having discussed the $\kappa\rightarrow\infty$ boundary actions for the family of fundamental Higgs $\mathrm{SU}(N)$ models we are in a position to say something about the phase transitions at the boundary. As the bulk is gapped for finite $\kappa$ we should expect that the boundary theory has a finite correlation length into the bulk making it a quasi-($\cramped{D-1}$)-dimensional boundary theory so that statements at $\kappa\rightarrow\infty$ hold also for finite $\kappa$.
@@ 3.5 Boundary Phase Transition Order and Universality Class
3.5.1 Group-Valued Higgs
The boundary degrees of freedom, Eq.˜59, are gauge-invariant composites transforming under the global $\mathcal{G}_{L}\times\mathcal{G}_{R}$ symmetry. Considering the case $\mathcal{G}=\mathrm{SU}(N)$, the coarse-grained (generally complex) matrix-valued fields are denoted $X^{ab}_{i}$ [48], where $a,b$ are color indices, which transform as $\mathbf{X}\rightarrow\mathbf{g}_{L}\mathbf{X}\mathbf{g}_{R}^{\dagger}$ under the symmetry. The Landau theory is
@@ 3.5.1 Group-Valued Higgs
Chiral models on a lattice, such as the one in Eq.˜60, have been studied for many years especially in two dimensions at large $N$, where they are integrable [49]. If one is interested in boundaries of four dimensional gauge theories, the three dimensional analogs of such models are of interest. One early work on $\mathcal{G}=SU(N)$ in three dimensions, Ref. [48], contains Monte Carlo results for $N=2,3,4$. For $N=2$ (Section˜3.2) the symmetry group is $\cramped{\mathrm{SU}(2)\times\mathrm{SU}(2)\simeq\mathrm{O}(4)}$ and the transition is continuous, consistent with $\mathrm{O}(4)$ criticality. For $N=3,4$ the numerical results reveal the transition to be first order in agreement with the mean field theory predictions.
3.5.2 Fundamental Higgs
The boundary model in the case of a fundamental Higgs has gauge-invariant degrees of freedom of the form $\Phi^{a}_{i}\equiv U_{i}^{ab}\phi_{i}^{b}$ where $a,b$ are the color indices, Eq.˜73. There is a single global $\mathcal{G}$ symmetry that acts from the left $\Phi\rightarrow g\Phi$. This will be broken spontaneously for sufficiently large $\beta$ (modulo Mermin-Wagner restrictions). The coarse-grained version of $\Phi^{a}$ is denoted $\Psi^{a}$ and the Landau theory for $\mathrm{SU}(N)$ is
which is invariant under $\mathrm{U}(N)$ transformations. This may instead be viewed as a theory of $2N$ real variables invariant under the enlarged $\mathrm{O}(2N)$ symmetry group. This Landau theory gives the impression that, for the three dimensional boundary of a four dimensional $\mathrm{SU}(N)$ gauge theory, there is a phase transition in the $\mathrm{O}(2N)$ universality class. In the case of $\mathrm{SU}(2)$ this is $\mathrm{O}(4)$ criticality as shown above at the level of the microscopic model both analytically and numerically. Note that the coupling of the $\Phi_{i}$ rotors in Eq.˜74 is invariant under $\mathrm{O}(2N)$ rotations, even though the microscopic action manifestly only has a global $\mathrm{SU}(N)$ symmetry.
It may seem surprising that the $\mathrm{SU}(N)$ invariant model exhibits $\mathrm{O}(2N)$ criticality. One simple check is that $\mathrm{SU}(N)$ has $N^{2}-1$ generators and that $\mathrm{SU}(N)\rightarrow\mathrm{SU}(N-1)$ symmetry breaking therefore has 2N12N-1 broken generators (corresponding to the number of Goldstone modes) which matches the count for $\mathrm{O}(2N)\rightarrow\mathrm{O}(2N-1)$ symmetry breaking where O($N$) has N(N1)/2N(N-1)/2 generators. But what about the terms that are $\mathrm{SU}(N)$ invariant but not $\mathrm{O}(2N)$ invariant? These terms are certainly present. An example is given by taking an operator $\Xi_{ab}\equiv\Psi_{a}\Psi^{\dagger}_{b}$ which transforms as $g\Xi g^{\dagger}$ and considering its determinant. This is only $\mathrm{U}(N)$ invariant, but it is also an irrelevant operator for $N>2$ in $D=4$. More precisely, the couplings in the action originating from the determinant have mass dimension $g_{\rm det}]=D+N(2-D)$ which, in dimensions higher than two, is negative for all but $N=2,3$ in three dimensions and $N=2$ in four dimensions. As we have seen, the case of $\mathrm{SU}(2)$ is special because it is identical to a problem with a group valued Higgs. So we have treated it separately. Therefore the remaining puzzle relates to $\mathrm{SU}(3)$ in three dimensions where the determinant coupling goes like $|\Psi^{\dagger}\Psi|^{3}$ and is marginal by power counting. As with the problem of scalar field theory in three dimensions [50], among other cases, we expect this sixth order term to be marginally irrelevant. Taking this together with our results for $\mathrm{SU}(2)$ we surmise that the $\mathrm{SU}(N)$ boundary theory phase transition is in the $\mathrm{O}(2N)$ universality class since terms breaking $\mathrm{O}(2N)$ down to $\mathrm{SU}(N)$ are irrelevant or marginally irrelevant.
3.5.3 Fate of the Un-Higgsed Subgroup: SU(3) Numerical Results
The discussion above for the fundamental Higgs assumed that the relevant critical degrees of freedom on the boundary are the $\Phi_{i}$, Eq.˜73. One may be concerned, however, about the residual fluctuations from the un-Higgsed $\mathrm{SU}(N-1)$ part of the gauge group. In particular, the $\Phi_{i}$ are composite degrees of freedom between a boundary link and its attached Higgs field, and that the boundary link also appears in the second term in the boundary action, Eq.˜74, a three-leg Wilson loop with projectors at the two bulk sites. However, since this term is not charged under the global $\mathrm{SU}(N)$ symmetry acting on the boundary links from the outside, we expect it to be irrelevant to the boundary criticality.
To verify this and our prediction of $\mathrm{O}(2N)$ criticality, we have performed simulations of 4D $\mathrm{SU}(3)$ gauge theory with fundamental Higgs. The Higgs field $\phi_{i}$ is a $3$-component complex unit vector at each site, and we measure the average of the Higgs field measured from the vacuum end of the boundary links, Eq.˜73, along with the boundary Wilson loops,
$\Phi$ $=\frac{1}{L^{3}}\sum_{i\in\partial X}(U_{i}\phi_{i})\,\in\,\mathbbm{C}^{3}$ (79)
[…]
The numerical results for the average Higgs field are presented in Fig.˜9 for $\kappa=3.0$. Figure˜9(a) shows its behavior for various system sizes, while Fig.˜9(b) shows the behavior of its variance. The behavior is consistent with that of an order parameter at a second order phase transition, approaching zero at small $\beta$ and continuously increasing from $\beta>\beta_{c}\approx 4.19$, with a diverging susceptibility. Figure˜9 shows the Binder cumulant scaled by the 3D $\mathrm{O}(6)$ critical exponent $\nu\approx 0.789$ [51], which shows a clear collapse to a universal scaling function. Furthermore, Fig.˜10 shows the behavior of $W_{P}$, which shows no system-size dependence and smooth behavior, verifying that this term is irrelevant at the transition. We therefore conclude that the $\mathrm{SU}(3)$ Higgs model in $D=4$ indeed demonstrates a boundary phase transition in the $3D$ $\mathrm{O}(6)$ universality class, even though a single fundamental Higgs field does not remove all bulk degrees of freedom when $\kappa\to\infty$, in agreement to the general picture of boundary $\mathrm{O}(2N)$ criticality on the boundary of $\mathrm{SU}(N)$ fundamental Higgs models.
@@ 3.5.3 Fate of the Un-Higgsed Subgroup: SU(3) Numerical Results
Figure 10: The expectation and variance of the boundary $\mathrm{SU}(N-1)$ Wilson loop, Eq.˜80, which appears as one of the two terms in the $\cramped{\kappa\to\infty}$ limit of the fundamental-Higgs $\mathrm{SU}(N)$ boundary action in Eq.˜74. This quantity shows no system-size dependence, indicating it plays no role in the boundary phase transition (cf. the boundary order parameter in Fig.˜9).
3.6 Summary and Discussion: Non-Abelian Case